# Improving Signed Propagation of Graph Neural Network Under Multiple Classes

### v6

@@ V-A1 Edge weight calibration for GCN
$\sum_{j\in\mathcal{N}_{i}}\left({k(1-e)+ke\times c\over d_{ij}}b_{i}+{k^{\prime}(1-e)\times c+k^{\prime}e\over d_{ij}}(1-b_{i})\right)$
$={k\over d_{i}+1}+{\{(c-1)ke+k\}b_{i}+(1-c)k^{\prime}e(1-b_{i})\over d_{i}+1}d^{\prime}$ (32)
@@
$Z$ $=kb_{i}+k^{\prime}(1-b_{i})-\{(c-1)ke+k\}b_{i}-(1-c)k^{\prime}e(1-b_{i})$
$=(1-c)keb_{i}+\{1+(c-1)e\}k^{\prime}(1-b_{i})$ (33)
@@
$\begin{gathered}Z\in\begin{cases}e+b_{i}-1-ce&\text{if}\,\,\cos(k^{\prime},k)=-1\\ -(e+b_{i}-1-ce)-2eb_{i}+2eb_{i}c&\text{if}\,\,\cos(k^{\prime},k)=1,\end{cases}\end{gathered}$ (34)
@@
V-A2 Edge weight calibration for GAT
Since the attention value of GAT is always positive, we can induce a similar equation as above. Thus, instead of similar notational derivation, we show that the attention mechanism is not sufficient to capture the angular similarity of two nodes. Generally, the formulation of attention is defined as follows:
$a_{ij}=\vec{a}^{T}[W\vec{h}_{i}||W\vec{h}_{j}]$ (35)
The equation above implies that the attention weight is determined by the summation of distances between the attention vector $\vec{a}=[\vec{a}_{1}||\vec{a}_{2}]$ and two nodes $W\vec{h}_{i}$, $W\vec{h}j$ since $a_{ij}=\vec{a}_{1}^{T}\cdot W\vec{h}_{i}+\vec{a}_{2}^{T}\cdot W\vec{h}_{j}$. Now, let’s assume that two neighbors, $j$ and $k$, are equidistant from the attention vector. In such cases, the attention weight assigned to both $j$ and $k$ towards the ego node will be identical, regardless of their directional positioning. Consequently, attention may not effectively capture cosine similarity, where we show that edge weight calibration can enhance positive GNNs by incorporating angular information.
Table I: (Q1) Node classification accuracy (%) with standard deviation on the six test datasets. Bold and underline indicates the $1^{st}$ and $2^{nd}$ best performances. Values in brackets represent the dissonance (Eq. 38) and methods with calibrations are marked with $\ddagger$.
  Datasets Cora Citeseer Pubmed Actor Chameleon Squirrel $\mathcal{H}_{g}$ (Eq. 1) 0.81 0.74 0.8 0.22 0.23 0.22   GCN 79.0 ± ${}_{\,\pm\,0.6}$ (0.17) 67.5 ± ${}_{\,\pm\,0.8}$ (0.29) 77.6 ± ${}_{\,\pm\,0.2}$ (0.53) 20.2 ± ${}_{\,\pm\,0.4}$ (0.29) 49.3 ± ${}_{\,\pm\,0.5}$ (0.19) 30.7 ± ${}_{\,\pm\,0.7}$ (0.31) GCN 81.0 ± ${}_{\,\pm\,0.9}$ (0.12) 71.3 ± ${}_{\,\pm\,1.2}$ (0.14) 77.8 ± ${}_{\,\pm\,0.4}$ (0.38) 21.7 ± ${}_{\,\pm\,0.6}$ (0.62) 49.4 ± ${}_{\,\pm\,0.6}$ (0.25) 31.5 ± ${}_{\,\pm\,0.6}$ (0.58) GAT 80.1 ± ${}_{\,\pm\,0.6}$ (0.22) 68.0 ± ${}_{\,\pm\,0.7}$ (0.25) 78.0 ± ${}_{\,\pm\,0.4}$ (0.45) 22.5 ± ${}_{\,\pm\,0.3}$ (0.28) 47.9 ± ${}_{\,\pm\,0.8}$ (0.17) 30.8 ± ${}_{\,\pm\,0.9}$ (0.27) GAT 81.4 ± ${}_{\,\pm\,0.4}$ (0.12) 72.2 ± ${}_{\,\pm\,0.6}$ (0.08) 78.3 ± ${}_{\,\pm\,0.3}$ (0.39) 23.2 ± ${}_{\,\pm\,1.8}$ (0.43) 49.2 ± ${}_{\,\pm\,0.4}$ (0.16) 30.3 ± ${}_{\,\pm\,0.8}$ (0.40) GIN 77.3 ± ${}_{\,\pm\,0.8}$ (0.33) 66.1 ± ${}_{\,\pm\,0.6}$ (0.29) 77.1 ± ${}_{\,\pm\,0.7}$ (0.47) 24.6 ± ${}_{\,\pm\,0.8}$ (0.51) 49.1 ± ${}_{\,\pm\,0.7}$ (0.26) 28.4 ± ${}_{\,\pm\,2.2}$ (0.48) APPNP 81.3 ± ${}_{\,\pm\,0.5}$ (0.15) 68.9 ± ${}_{\,\pm\,0.3}$ (0.21) 79.0 ± ${}_{\,\pm\,0.3}$ (0.42) 23.8 ± ${}_{\,\pm\,0.3}$ (0.49) 48.0 ± ${}_{\,\pm\,0.7}$ (0.34) 30.4 ± ${}_{\,\pm\,0.6}$ (0.69) GCNII 81.1 ± ${}_{\,\pm\,0.7}$ (0.08) 68.5 ± ${}_{\,\pm\,1.4}$ (0.13) 78.5 ± ${}_{\,\pm\,0.4}$ (0.20) 25.9 ± ${}_{\,\pm\,1.2}$ (0.43) 48.1 ± ${}_{\,\pm\,0.7}$ (0.21) 29.1 ± ${}_{\,\pm\,0.9}$ (0.24) H2GCN 80.6 ± ${}_{\,\pm\,0.6}$ (0.16) 68.2 ± ${}_{\,\pm\,0.7}$ (0.22) 78.5 ± ${}_{\,\pm\,0.3}$ (0.29) 25.6 ± ${}_{\,\pm\,1.0}$ (0.34) 47.3 ± ${}_{\,\pm\,0.8}$ (0.19) 31.3 ± ${}_{\,\pm\,0.7}$ (0.62) ACM-GCN 80.2 ± ${}_{\,\pm\,0.8}$ (0.24) 68.3 ± ${}_{\,\pm\,1.1}$ (0.17) 78.1 ± ${}_{\,\pm\,0.5}$ (0.31) 24.9 ± ${}_{\,\pm\,2.0}$ (0.46) 49.5 ± ${}_{\,\pm\,0.7}$ (0.20) 31.6 ± ${}_{\,\pm\,0.4}$ (0.54) HOG-GCN 79.7 ± ${}_{\,\pm\,0.4}$ (0.31) 68.2 ± ${}_{\,\pm\,0.6}$ (0.24) 78.0 ± ${}_{\,\pm\,0.2}$ (0.29) 21.5 ± ${}_{\,\pm\,0.5}$ (0.37) 47.7 ± ${}_{\,\pm\,0.5}$ (0.32) 30.1 ± ${}_{\,\pm\,0.4}$ (0.51) JacobiConv 81.9 ± ${}_{\,\pm\,0.6}$ (0.26) 69.6 ± ${}_{\,\pm\,0.8}$ (0.19) 78.5 ± ${}_{\,\pm\,0.4}$ (0.24) 25.7 ± ${}_{\,\pm\,1.2}$ (0.30) 52.8 ± ${}_{\,\pm\,0.9}$ (0.21) 32.0 ± ${}_{\,\pm\,0.6}$ (0.37) GloGNN 82.4 ± ${}_{\,\pm\,0.3}$ (0.33) 70.3 ± ${}_{\,\pm\,0.5}$ (0.35) 79.3 ± ${}_{\,\pm\,0.2}$ (0.26) 26.6 ± ${}_{\,\pm\,0.7}$ (0.38) 48.2 ± ${}_{\,\pm\,0.3}$ (0.27) 28.8 ± ${}_{\,\pm\,0.8}$ (0.43)   PTDNet 81.2 ± ${}_{\,\pm\,0.9}$ (0.24) 69.5 ± ${}_{\,\pm\,1.2}$ (0.42) 78.8 ± ${}_{\,\pm\,0.5}$ (0.44) 21.5 ± ${}_{\,\pm\,0.6}$ (0.33) 50.6 ± ${}_{\,\pm\,0.9}$ (0.17) 32.1 ± ${}_{\,\pm\,0.7}$ (0.34) PTDNet 81.9 ± ${}_{\,\pm\,0.6}$ (0.20) 71.1 ± ${}_{\,\pm\,0.8}$ (0.31) 79.0 ± ${}_{\,\pm\,0.2}$ (0.38) 22.7 ± ${}_{\,\pm\,0.6}$ (0.19) 50.9 ± ${}_{\,\pm\,0.3}$ (0.15) 32.3 ± ${}_{\,\pm\,0.5}$ (0.30)   GPRGNN 82.2 ± ${}_{\,\pm\,0.4}$ (0.25) 70.4 ± ${}_{\,\pm\,0.8}$ (0.43) 79.1 ± ${}_{\,\pm\,0.1}$ (0.26) 25.4 ± ${}_{\,\pm\,0.5}$ (0.55) 49.1 ± ${}_{\,\pm\,0.7}$ (0.25) 30.5 ± ${}_{\,\pm\,0.6}$ (0.36) GPRGNN 84.7 ± ${}_{\,\pm\,0.2}$ (0.04) 73.3 ± ${}_{\,\pm\,0.5}$ (0.05) 80.2 ± ${}_{\,\pm\,0.2}$ (0.11) 28.1 ± ${}_{\,\pm\,1.3}$ (0.33) 51.0 ± ${}_{\,\pm\,0.4}$ (0.18) 31.8 ± ${}_{\,\pm\,0.4}$ (0.16) FAGCN 81.9 ± ${}_{\,\pm\,0.5}$ (0.15) 70.8 ± ${}_{\,\pm\,0.6}$ (0.17) 79.0 ± ${}_{\,\pm\,0.5}$ (0.31) 25.2 ± ${}_{\,\pm\,0.8}$ (0.66) 46.5 ± ${}_{\,\pm\,1.1}$ (0.25) 30.4 ± ${}_{\,\pm\,0.4}$ (0.64) FAGCN 84.1 ± ${}_{\,\pm\,0.4}$ (0.09) 73.8 ± ${}_{\,\pm\,0.5}$ (0.08) 79.7 ± ${}_{\,\pm\,0.2}$ (0.16) 27.6 ± ${}_{\,\pm\,0.5}$ (0.42) 48.8 ± ${}_{\,\pm\,0.7}$ (0.13) 31.3 ± ${}_{\,\pm\,0.5}$ (0.37) GGCN 81.0 ± ${}_{\,\pm\,1.2}$ (0.38) 70.7 ± ${}_{\,\pm\,1.6}$ (0.30) 78.2 ± ${}_{\,\pm\,0.4}$ (0.47) 22.5 ± ${}_{\,\pm\,0.5}$ (0.47) 48.5 ± ${}_{\,\pm\,0.7}$ (0.15) 30.2 ± ${}_{\,\pm\,0.7}$ (0.40) GGCN 83.9 ± ${}_{\,\pm\,0.8}$ (0.07) 73.0 ± ${}_{\,\pm\,0.4}$ (0.05) 78.9 ± ${}_{\,\pm\,0.3}$ (0.29) 24.6 ± ${}_{\,\pm\,0.4}$ (0.26) 50.0 ± ${}_{\,\pm\,0.4}$ (0.07) 31.1 ± ${}_{\,\pm\,0.6}$ (0.15)
Table II: Statistical details of the six benchmark datasets
V-B (Parameter Update) Confidence Calibration
Theorem 4.6 shows that signed messages increase the uncertainty of predictions. To redeem this, we propose the calibration of predictions. This scheme boasts several advantages such as being free from path configuration, cost-efficient, and powerful. Here, we claim that various loss functions can be used for confidence calibration as below:
$\begin{gathered}\mathcal{L}_{conf}\in\begin{cases}{1\over N}\sum_{i=1}^{N}(-\text{max}(\widehat{y}_{i})+\text{submax}(\widehat{y}_{i}))\\ \sum_{i=1}^{N}||I-\widehat{y}_{i}\widehat{y}_{i}^{T}||\end{cases}\end{gathered}$ (36)
The first one suppresses the maximal values and reduces the differences between the maximal and sub-maximal values, while the second one is a well-known orthogonal constraint. Both of them can be used to relieve the conflict evidence of predictions. Please note that any penalization strategy reducing entropy can be applied as a method of confidence calibration. One may argue that this is quite similar to the prior scheme [16, 41], but we provide a new understanding that signed propagation may incur higher uncertainty (Thm. 4.6).
@@ V-A2 Edge weight calibration for GAT
Algorithm 1 Pseudo-code: FAGCN [11] with our method

### v7

@@ E. Edge Weight and Confidence Calibration
(86) $\mathbb{E}(h^{(1)}_{i}|v_{i},d_{i})$ $={k\over d_{i}+1}+\sum_{j\in\mathcal{N}_{i}}\left({k(1-e)+ke\times s\over\sqrt{(d_{i}+1)(d_{j}+1)}}b_{i}+{k^{\prime}(1-e)\times s+k^{\prime}e\over\sqrt{(d_{i}+1)(d_{j}+1)}}(1-b_{i})\right)$
(87) $={k\over d_{i}+1}+\sum_{j\in\mathcal{N}_{i}}\left({\{(s-1)ke+k\}b_{i}+(1-s)k^{\prime}e(1-b_{i})\over\sqrt{(d_{i}+1)(d_{j}+1)}}\right)$
(88) $={k\over d_{i}+1}+{\{(s-1)ke+k\}b_{i}+(1-s)k^{\prime}e(1-b_{i})\over d_{i}+1}d^{\prime}$
@@
(89) $Z$ $=kb_{i}+k^{\prime}(1-b_{i})-\{(s-1)ke+k\}b_{i}-(1-s)k^{\prime}e(1-b_{i})$
(90) $=(1-s)keb+\{1+(s-1)e\}k^{\prime}(1-b_{i})$
@@
(91) $\begin{gathered}Z\in\begin{cases}e+b_{i}-1-se,&\text{if}\,\,\cos(k^{\prime},k)=-1\\ e+b_{i}-1-se+2eb_{i}(s-1),&\text{if}\,\,\cos(k^{\prime},k)=1,\end{cases}\end{gathered}$
@@
Figure 8. Attention Mechanism
E-1. Edge weight calibration for GAT (Velickovic et al., 2017). Now, we show that the attention mechanism is not sufficient to capture the similarity of two nodes. Generally, the formulation of attention is defined as below:
(92) $a_{ij}=\vec{a}^{T}[W\vec{h}_{i}||W\vec{h}_{j}],$
which is the summation of distances between the attention vector $\vec{a}$ and two nodes $W\vec{h}_{i},W\vec{h}_{j}$ as follows: $a_{ij}=\vec{a}_{1}^{T}\cdot W\vec{h}_{i}+\vec{a}_{2}^{T}\cdot W\vec{h}_{j}$ since attention vector can be decomposed into two vectors $\vec{a}=[\vec{a}_{1}||\vec{a}_{2}]$. Let us assume two neighbors $j$ and $k$ have the same distance from the attention vector. In this case, we can easily infer that the neighbors have identical attention weights since $d+d_{1}=d+d_{2}$ and the normalization would remain the same. However, their cosine similarity differs, and we can observe that the attention mechanism may fail to capture this information.
E-2. Confidence calibration and uncertainty.
For brevity, assume a star-like subgraph $\mathcal{G}$ consists of an ego node ($i$) and its neighboring nodes ($j$) with homophily ($\mathcal{H}_{\mathcal{G}}$). Also, let us assume that the label of an ego is 0 ($y_{i}=0$) under 3 classes $|C|=3$, which is the only node used for training. Firstly, we can easily infer that the prediction of an ego will converge as $\widehat{y}_{i}=[1,0,0]$. Conversely, the neighboring nodes will either be $\widehat{y}_{i}=[0,0.5,0.5]$ or $\widehat{y}_{i}=[1,0,0]$ depending on the sign of connected edges. In this case, the Shannon Entropy can be calculated as $H([1,0,0])=0$ and $H([0,0.5,0.5])=\log(c-1)=1$. Thus, we can infer that $\mathbb{E}[\mathcal{G}]\approx\mathcal{H}_{\mathcal{G}}\cdot\log(c-1)$ considering that the neighboring nodes are independent and identically distributed. Now, let’s apply the confidence calibration as below:
(93) $\mathcal{L}_{conf}={1\over N}\sum_{i=1}^{N}(-\text{max}(\widehat{y}_{i})+\text{submax}(\widehat{y}_{i}))$
Firstly, under perfect homophily $\mathcal{H}_{\mathcal{G}}=1$, the entropy of all neighbors $H([1,0,0])$ that are connected with positive edges will remain the same as 0, where we can induce that confidence calibration is unnecessary $H(\mathbb{E}[\widehat{y}_{cal}])=H(\mathbb{E}[\widehat{y}_{s}])$ under fully converged condition. However, this condition may not hold under real-world benchmark graphs. Conversely, in case $\mathcal{H}_{\mathcal{G}}\neq 1$, neighbors with signed edges will be $H([0,0.5+\lambda,0.5-\lambda])$ as we explained in Eq. 74. Since $\alpha$ is proportional to the $\lambda\mathcal{L}_{conf}$ (Eq. 32), naively assuming that $10\leq\lambda<$ 1, the entropy decreases as $H([0,0.6,0.4])=0.971$. As such, the entropy after confidence calibration follows:
(94) $\mathcal{H}_{\mathcal{G}}\cdot\{{1\over c-1}\log({c-1\over 1+\lambda})+{c-2\over c-1}\log({c-2\over 1-\lambda})\}\leq\mathcal{H}_{\mathcal{G}}\cdot\{{1\over c-1}\log({c-1\over 1+\lambda})+{c-2\over c-1}\log({c-1\over 1-\lambda})\}\leq\mathcal{H}_{\mathcal{G}}\cdot\log({c-1\over 1+\lambda})$
Given that the entropy of using signed messages only is $H(\mathbb{E}[\widehat{y}^{t}_{s}])$, the expectation of converged entropy ($t\rightarrow\infty$) after confidence calibration ($\mathbb{E}[\widehat{y}^{t}_{cal}]$) is given by:
[…]
which completes the proof ∎.
In summary, confidence calibration reduces the entropy of positive message-passing. However, it significantly reduces uncertainty in signed propagation as the homophily of the graph decreases. In this context, the frequency of conflict predictions is proportional to the heterophily ratio, which typically exhibits the highest entropy.
F. Pseudo-code and Time Complexity of Our Algorithm
@@
F-1. The pseudo-code of calibrated FAGCN (Bo et al., 2021) is as below. Instead of FAGCN, any GNNs can be integrated with our method
Algorithm 1 Pseudo-code: FAGCN (Bo et al., 2021) with our method